Calculate the molar mass, and give the mass % of each element in hydrated nickel(II) chloride, NiCl2 6H20.
Answer #1

1)

Molar mass of NiCl2.6H2O,
MM = 1*MM(Ni) + 2*MM(Cl) + 12*MM(H) + 6*MM(O)
= 1*58.69 + 2*35.45 + 12*1.008 + 6*16.0
= 237.686 g/mol
Answer: 237.686 g/mol


2)
Lets calculate the mass of each element
mass of Ni = 1*MM(Ni)
= 58.69 g


mass of Cl = 2*MM(Cl)
= 70.9 g


mass of H = 12*MM(H)
= 12.096 g


mass of O = 6*MM(O)
= 96 g



mass % = (mass of element *100)/molar mass of compound

Now calculate mass percent of each element
Mass % of Ni = (58.69*100)/237.686
= 24.6922 %
Mass % of Cl = (70.9*100)/237.686
= 29.8293 %
Mass % of H = (12.096*100)/237.686
= 5.0891 %
Mass % of O = (96*100)/237.686
= 40.3894 %

Answer:
Mass % of Ni = 24.69 %
Mass % of Cl = 29.83 %
Mass % of H = 5.09 %
Mass % of O = 40.39 %